OpenStudy (anonymous):

Determine whether the function is even, odd, or neither f(x)=3x^3+5

OpenStudy (anonymous):

well, you have an exponent of 3, that means that with -4 you will not get the same value as with positive 4. therefore it is odd.

OpenStudy (anonymous):

f(x)=3x^3+5 f(x)= 3(4)^3+5 = 192+5 = 197 f(x)= 3(-4)^3+5 = -192+5 = -187 if your function had an even exponent of 2, 4, and such then it would have been called an "even function" .